A sum of $2700 is to be given in the form of 63 prizes. If the prize is of either $100 or $25, find the number of prizes of each type.
Answer by stanbon(75887)

A sum of $2700 is to be given in the form of 63 prizes. If the prize is of either $100 or $25, find the number of prizes of each type.
----
Equations:
Quantity Eq: h + t = 63
Value Eq: 100h+25t = 2700
---------------------------------
Modify:
4h + 4t = 4*63
4h + t = 108
-----
Subtract and solve for "t":
3t = 144
t = 48 (# of $25 prizes)
-----
Solve for "h":
h + t = 63
h = 63-48
h = 15 (# of $100 prizes)
=============================
Cheers,
Stan H.
===============